OpenStudy (anonymous):

Heelp!! 3_Resolva equations. a) 2x + (4-x) = 3 - (6-3x) b) 4 (x-2) -2 (x-1) = 4 c) 10 - (x +2) = 9 + (x-5) d) 4-x (3x-6) +8 = x e) 3X-[2 - (1-x)] = 5x

OpenStudy (nincompoop):

let me show you how to solve one of them 2x + (4- x) = 3 it can be broken down to 2x + 4 - x = 3 combine all the x's 2x - x + 4 = 3 isolate x since it is what we are solving for 2x - x + 4 + (-4) = 3 + (-4) x = -1

OpenStudy (nincompoop):

2x + (4-x) = 3 - (6-3x) can be broken down to 2x + 4 - x = 3 - 6 + 3x do the same steps I did

OpenStudy (anonymous):

its 7/2 or 3.5

OpenStudy (anonymous):

HI :DDDD

OpenStudy (anonymous):

Hi!